OpenStudy (anonymous):

f(x)=sin2x g(x)=x^2-7 find g(f(x))

OpenStudy (anonymous):

To find g(f(x)) put sin2x in place of x in g(x).. Getting??

OpenStudy (anonymous):

and what would that become?

OpenStudy (anonymous):

You show me then I will tell you whether it is right or wrong..

OpenStudy (anonymous):

(sin2x)^2-7

OpenStudy (anonymous):

Yes now take the square of sin2x..

OpenStudy (anonymous):

you mean (sin2x)(sin2x)

OpenStudy (anonymous):

No.. x*x = x^2

OpenStudy (anonymous):

\[\huge (\sin2x)^2 = \sin^22x\]

OpenStudy (anonymous):

Getting??

OpenStudy (anonymous):

erm not that no.

OpenStudy (anonymous):

Are you provided with choices or not?

OpenStudy (anonymous):

\[\huge (\sin2x)^2 = \sin^22x\] These are one and the same thing..

OpenStudy (anonymous):

so how do you get from one to the other?

OpenStudy (anonymous):

See in case of trigonometry: \[\sin^2x\]\[\huge \sin^2x \implies (sinx)^2\]

OpenStudy (anonymous):

are you provided with the choices??

OpenStudy (anonymous):

you are right if that is what you mean.

OpenStudy (anonymous):

Yeah I mean it..

OpenStudy (anonymous):

but I do understand how sin^2 x is the same as (sin2x)^2